OpenStudy (anonymous):

What is the solution to the rational equation (3)/(2)-(1)/(x-5)=(-2)/(2x-10)? Can someone please help me!

OpenStudy (anonymous):

You can cross multiply to get your answer. Don't forget to check the answers you get to make sure that your answers work and aren't extraneous (A solution of a simplified version of an equation that does not satisfy the original equation.) http://www.mathwords.com/e/extraneous_solution.htm

OpenStudy (anonymous):

okay @wonder291

OpenStudy (anonymous):

but what exactly do i do to the other side of the equation

OpenStudy (anonymous):

\[\frac{ 3 }{ 2 }-\frac{ 1 }{ x-5 }=\frac{ -2 }{ 2x-10 }\] You could multiple the entire equation by the least common multiple of all the denominators... Then you could combine like terms

OpenStudy (anonymous):

could you guide me through the problem im struggling on this a lot.

OpenStudy (anonymous):

The Least Common Multiple (LCM) is basically the smallest (non-zero) number that is a multiple of two or more numbers. http://www.mathsisfun.com/least-common-multiple.html The same works for expressions with variables. For example, in this problem the numbers we have (to find the LCM of) are 2, x-5, and 2x-10.

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

Do you still need help with anything?

OpenStudy (anonymous):

Yes i am really trying my hardest to solve this problem but i just get confused on it .

OpenStudy (anonymous):

What step are you on now?

OpenStudy (anonymous):

cross multiplying

OpenStudy (anonymous):

Let me show you an example with other numbers \[\frac{ 2 }{ 3 }=\frac{ 4 }{ 6 }\] Cross multiplying in this example means you multiple one numerator by the other denominator.  2*6=3*4 12=12 The same would work with these expressions with variables.

OpenStudy (anonymous):

this is what i got 3x-15-1=(-2)/(2x-10)

OpenStudy (anonymous):

\[\frac{ 3 }{ 2 }-\frac{ 1 }{ x-5 }=\frac{ -2 }{ 2x-10 }\] What did you get for the least common multiple?

OpenStudy (anonymous):

i Really dont know :(

OpenStudy (anonymous):

It's ok, I'll help you... To find the least common multiple (denominator because they're a fraction) for the following, 2, x-5, and 2x-10. Try multiplying factors together so that you only have 1 factor. Try not to multiply all of them at once; try two at a time.

OpenStudy (anonymous):

okay i did it but it did not work

OpenStudy (anonymous):

another way to find it is to see if some factors divide into other factors

OpenStudy (anonymous):

for example, for 5, 2x+1, 10x+5 10x+5 is divisible by 5; and it is divisible by 2x+1; therefore it is the least common multiple

OpenStudy (anonymous):

It will work for 2, x-5, and 2x-10.

OpenStudy (anonymous):

for the example, 10x+5 is the least common multiple

OpenStudy (anonymous):

okay i multiplied them all by the lcm what do i do next

OpenStudy (anonymous):

You can combine like terms. for example: 2x+1+5x 2x and 5x are like terms. so you end up with 7x+1

OpenStudy (anonymous):

but before that, make sure you multiply everything out

OpenStudy (anonymous):

okay is the answer X=5

OpenStudy (anonymous):

substitute the answer you get back into the equation (to check), and remember that the denominator can never equal 0!

OpenStudy (anonymous):

i Believe this is right x=5

OpenStudy (anonymous):

If none of your solutions work, then there is no solution to the equation! (sometimes this happens)

OpenStudy (anonymous):

Oh so its no solution because it was'nt quite working

OpenStudy (anonymous):

\[\frac{ 3 }{ 2 }-\frac{ 1 }{ x-5 }=\frac{ -2 }{ 2x-10 }\] Yes, just for this rational equation.

OpenStudy (anonymous):

thankyou soo much for the guidance

OpenStudy (anonymous):

No problem! :)